# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v1

@@ 3.2 HGNN layer for particle reconstruction and pruning
HGNNs improve upon homogeneous GNNs by natively handling multiple node and edge types with type-specific representations and update functions, unlike the uniform approach of homogeneous GNNs. This inherent flexibility makes HGNNs better suited for modeling diverse systems, like physical interactions. While homogeneous GNNs can mimic this heterogeneity using techniques like one-hot encoding and padding, they lack the inherent inductive bias of HGNNs, often leading to suboptimal representations and making it difficult to set distinct learning objectives for different entity types.
@@
Figure 1: (a) Heterogeneous graph representation (below) and HGNN architecture (above) for a simultaneous beauty hadron reconstruction and PV association. In the graph solid lines indicate all true edges and their class labels, while dotted lines provide examples of false edges with label zero. (b) Heterogeneous modification in blue to the GNN layer updates from Battaglia et al[45] in red. The addition of pruning tasks is shown in light brown.
@@
Figure 1(a) illustrates the heterogeneous nature of a LHCb collision event with multiple reconstructed PVs, which produce $\mathcal{O}$(100-1,000) charged particle tracks. Data relations within the collision event can be naturally represented by a heterogeneous graph, with PV and track node types, and track-track and PV-track edges, which signify relations between the various node objects. Here, we chose not to include PV-PV edges as they are not directly associated with any learning objective. Our previous DFEI algorithm assigned each track to a PV with a minimum impact parameter and appended the corresponding PV’s coordinates as static track‑node features. This both capped performance, since the model couldn’t learn to select the correct PV dynamically, and increased memory requirements by duplicating the same coordinates across tracks. In contrast, a heterogeneous graph explicitly encodes PV-track edges, allowing the end‑to‑end learning of associations and eliminating repetitions of PV data.
@@
To enable learning graph representations for joint beauty hadron reconstruction and PV association, we propose a heterogeneous extension of the GNN layer of Battaglia et al.. We denote track and PV node representations, $v^{i}_{\rm tr}$ and $v^{j}_{\rm pv}$, track-track and PV-track edge representations, $e^{k}_{\rm tr}$ and $e^{l}_{\rm pv-tr}$. The corresponding edge, node and global representation updates for our HGNN layer are,
{split}$\split e^{\prime k}_{\rm tr}=\phi^{e_{\rm tr}}(e^{k}_{\rm tr},v^{r_{k}}_{\rm tr},v^{s_{k}}_{\rm tr},u)\qquad&\bar{e}^{\prime i}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm tr}\})\\ e^{\prime l}_{\rm pv-tr}=\phi^{e_{\rm pv-tr}}(e^{l}_{\rm pv-tr},v^{l}_{\rm tr},v^{l}_{\rm pv},u)\qquad\bar{e}^{\prime i}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm pv-tr}\})\\ v^{\prime i}_{\rm tr}=\phi^{v_{\rm tr}}(v^{i}_{\rm tr},\bar{e}^{\prime i}_{\rm tr},\bar{e}^{\prime i}_{\rm pv-tr},u)\qquad\bar{e}^{\prime j}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm pv}}(\{E^{\prime j}_{\rm pv-tr}\})\\ v^{\prime j}_{\rm pv}=\phi^{v_{\rm pv}}(v^{j}_{\rm pv},\bar{e}^{\prime j}_{\rm pv-tr},u)\qquad\bar{e}^{\prime}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow u}(\{E^{\prime}_{\rm tr}\}),\bar{v}^{\prime}_{\rm tr}=\rho^{v_{\rm tr}\rightarrow u}(\{V^{\prime}_{\rm tr}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime}_{\rm tr},\bar{e}^{\prime}_{\rm pv-tr},\bar{v}^{\prime}_{\rm tr},\bar{v}^{\prime}_{\rm pv},u)\qquad\bar{e}^{\prime}_{\rm tr-pv}=\rho^{e_{\rm tr-pv}\rightarrow u}(\{E^{\prime}_{\rm tr-pv}\}),\bar{v}^{\prime}_{\rm pv}=\rho^{v_{\rm pv}\rightarrow u}(\{V^{\prime}_{\rm pv}\})\;,$ (2)
@@
where the HGNN layer update begins with edge updates, $e^{\prime k}_{\rm tr}$ and $e^{\prime l}_{\rm pv-tr}$, applying MLPs $\phi^{e_{\rm tr}}$ and $\phi^{e_{\rm pv-tr}}$. Meanwhile, node updates, $v^{\prime i}_{\rm tr}$ and $v^{\prime j}_{\rm pv}$ with MLPs $\phi^{v_{\rm tr}}$ and $\phi^{v_{\rm pv}}$ subsequently take as input aggregated edge representations for multiple edge types. The inputs of the $v^{\prime i}_{\rm tr}$ update includes the aggregated track edges $\bar{e}^{\prime i}_{\rm tr}$ and PV-track edges $\bar{e}^{\prime i}_{\rm pv-tr}$. Meanwhile, the update $v^{\prime j}_{\rm pv}$ includes the aggregation of PV-track edges to PV nodes, $\bar{e}^{\prime j}_{\rm pv-tr}$. Lastly, the global update uses as input aggregations of all node and edge types globally, $\bar{e}^{\prime}_{\rm tr}$, $\bar{e}^{\prime}_{\rm pv-tr}$, $\bar{v}^{\prime}_{\rm tr}$, $\bar{v}^{\prime}_{\rm pv}$. Figure 1(b), demonstrates the various HGNN layer computations of equation 2 for node and edge inputs from tracks (red) and PVs (blue).
@@
To enhance the scalability of the HGNN in potential real-time data acquisition settings, we incorporate edge and node pruning directly into its layers. The goal is to learn node and edge scores that can be used to remove irrelevant nodes and edges at inference time. To this end, we introduce node and edge scores, $\hat{y}^{v}=\sigma(\psi^{v}(v^{\prime}))$ and $\hat{y}^{e}=\sigma(\psi^{e}(e^{\prime}))$, where $\psi^{v}$ and $\psi^{e}$ are MLPs operating on the updated node and edge representations $v^{\prime}$ and $e^{\prime}$ from Equation 2, and $\sigma$ denotes the Sigmoid function. These probability scores can be trained in two ways: using a Binary Cross-Entropy loss $\mathcal{L}_{\rm BCE}(\hat{y}^{v/e},y^{v/e})$ when domain knowledge provides suitable ground-truth pruning labels $y^{v}$ and $y^{e}$, or implicitly, by using the predicted scores $\hat{y}^{v}$ and $\hat{y}^{e}$ as weights within the HGNN aggregation functions. We explore both approaches in our ablation studies in Section 4.1.
@@
Pruning can be performed either via a top-$k$ operation or by introducing a threshold $y_{\rm cut}$. In this paper, we adopt the latter approach given better observed performance and greater interpretability in terms of purity. Furthermore, we focus on the application of pruning during inference to boost speed and remove redundant background edges and nodes to enhance reconstruction performance. We acknowledge that integrating pruning into training (i.e., ‘pruning-aware’ training, a form of dynamic graph construction) could further improve reconstruction. Despite not employing pruning-aware training, our empirical results show that incorporating extra MLPs per node and edge improves the algorithm’s performance, as detailed in subsequent sections.
@@ 3.3 HGNN architecture and objective loss
The full HGNN architecture is shown in Figure 1(a), which processes as input $G_{\rm in}$ including track and PV node and edge input features through an encoder layer, several HGNN layers with integrated graph pruning and finally a decoder layer resulting in $G_{\rm out}$. The encoder and decoder layers consist of MLP node, edge and global updates, $\chi(v)$, $\chi(e)$ and $\chi(u)$. The encoded graph representation, $G_{\rm enc}$, is concatenated with the output graph representation from each HGNN layer to preserve important information from the input features and prevent oversmoothing [46]. Hyperparameters of the HGNN include the number of HGNN layers ($n^{\rm HGNN}_{\rm layers}=8$) and layers within MLPs ($n^{\rm MLP}_{\rm layers}=4$), the MLP hidden channel dimensions ($d^{\chi}_{\rm h}=d^{\phi}_{\rm h}=128$ and $d^{\psi}_{\rm h}=16$) and the MLP output dimensions ($d^{\chi}_{\rm h}=d^{\phi}_{\rm h}=16$ and $d^{\psi}_{\rm h}=1$). All MLPs employ ReLU activation functions and batch norm, meanwhile aggregation functions use summation. The hyperparameters were chosen on the basis of an observed saturation in the validation loss with model complexity.
The HGNN is trained to accomplish beauty hadron reconstruction, graph pruning and PV association tasks by minimising a multi-objective loss,
@@
where in the first term $\mathcal{L}_{\rm CE}$ denotes a class weighted cross-entropy loss, $e_{\rm out}$ is the output HGNN edge representation and $y^{\rm LCA}$ is a multi-class target for the LCA reconstruction task. The subsequent terms represent the edge and node pruning tasks introduced in Section 3.2, where the significance of the corresponding pruning task relative to the LCA task is scaled with $\beta$ parameters. The pv-tr edge pruning task effectively performs a PV association by selecting the pv-tr edge with the highest probability score $\hat{y}^{e_{\rm pv-tr}}$ for a given track. By default trainings used $\beta^{e_{\rm tr}}=30$, $\beta^{v_{\rm tr}}=3$ and $\beta^{e_{\rm pv-tr}}=1$, which were not optimal, therefore the sensitivity of the multi-objective optimisation to these parameters is further investigated in  B.
We benchmark the HGNN architecture against a GNN, using GNN layer updates according to equation 1, and an equivalent pruning mechanism. The GNN only considers tracks and their interconnections as nodes and edges, respectively. Meanwhile, it uses the same objective loss and hyperparameter values but with the removal of the pv-tr edge pruning task. For both GNN and HGNN architectures, we also quantify the inclusion of weighted aggregations during message passing and denote the corresponding architectures as WGNN and WHGNN.
@@ 3.4 Datasets and training procedure
The model is trained, validated and tested with the publicly available datasets from our previous publication DFEI [47]. The datasets were produced with a custom simulation environment using PYTHIA8 [48] and EvtGen [49] to simulate particle collision conditions anticipated for LHCb Run 3. Only events containing at least one beauty hadron are considered; these hadrons are then decayed with EvtGen into a variety of known modes. All results in this paper focus exclusively on charged particles produced within the LHCb geometrical acceptance and the Vertex Locator region. Particles outside these regions and neutral particles are not considered, meaning that they are also excluded from the ground truth heavy-hadron decay chains. Additional exclusive datasets, in which one of the beauty hadrons are required to decay to several exclusive decays, are used to further evaluate the model performance.
@@
As depicted in Figure 1, the collision events are represented as a heterogeneous graph. Each node and edge type has an input feature representation. For charged particles (tracks), the input feature representations include the track origin point, $(x_{\rm tr},y_{\rm tr},z_{\rm tr})$, the track momenta ($p_{x},p_{y},p_{z}$) and track charge $q$. Meanwhile, PV nodes are represented by their position coordinate, $(x_{\rm pv},y_{\rm pv},z_{\rm pv})$. In contrast, for GNN benchmarks, we adopt the DFEI’s approach of appending the coordinates of the PV with minimum impact parameter to the features of each track. Although highly discriminating timing information is anticipated for PV association in LHCb Upgrade II, this study limits its scope to using only these defined positional and kinematic features.
@@
Initially, a fully connected graph is defined between all track and PV nodes with track-track and track-PV edges. The input edge representation for PV-track edges includes solely the impact parameter of the track with respect to the PV. Meanwhile, for track-track edges features, include the angle $\theta$ between the three-momentum directions of the two particles, the difference $\Delta z_{\rm tr}$ in origin $z_{\rm tr}$, a boolean indicating a shared PV according to minimum impact parameter and the momentum-transverse distance, which is the distance between the origin points in a plane transverse to the momentum direction. We adopted a loose prefiltering of track-track edges with 99% efficiency from the DFEI [22], which requires edges to satisfy $\theta<0.26$ rad. or the condition of a shared PV according to minimum impact parameter. For the remaining edges, the LCAG ground-truth is determined $y^{\rm LCA}$, where $y^{\rm LCA}=0$ indicates no shared ancestors. The non-zero values are restricted to $y^{\rm LCA}=1,2,3$, which were found to be sufficient to represent the decay hierarchies present in the simulation dataset.
@@
The training and validation datasets consist of 40,000 and 10,000 events respectively, which are equivalent to those used in our previous work. Nominal training for the ablation studies was carried out on an NVIDIA L40S GPU with a batch size of 12. For ablation studies, a base schedule consisted of 30 epochs at a learning rate of $\alpha=10^{-3}$, followed by two epochs at $\alpha=10^{-4}$. Before evaluating the reconstruction performance of selected architectures, we fine-tuned with two additional epochs at $\alpha=5\times 10^{-5}$ and two at $\alpha=2\times 10^{-5}$. During all training phases, training and validation losses as well as classification accuracies were monitored (see  A for training and validation loss curves). The final performance was measured on a test set of 10,000 independent inclusive beauty-hadron decay events and on multiple exclusive decay datasets, each containing 5,000 events.

### v2

@@ 3.2 HGNN layer for particle reconstruction and pruning
Unlike the uniform approach of homogeneous GNNs, HGNNs improve upon homogeneous GNNs by natively handling multiple node and edge types with type-specific representations and update functions. This inherent flexibility makes HGNNs better suited for modeling diverse systems such as physical interactions. Although homogeneous GNNs can mimic this heterogeneity using techniques such as one-hot encoding and padding, they lack the inherent inductive bias of HGNNs, often leading to suboptimal representations and making it difficult to set distinct learning objectives for different entity types.
@@
Figure 1: (a) Heterogeneous graph representation (above) and HGNN architecture (below) for a simultaneous beauty hadron reconstruction and PV association. In the graph solid lines indicate all true edges and their class labels, while dotted lines provide examples of false edges with label zero. (b) Heterogeneous modification in blue to the GNN layer updates from Battaglia et al[45] in red. The addition of pruning tasks is shown in light brown.
@@
Figure 1(a) illustrates the heterogeneous nature of a LHCb collision event with multiple reconstructed PVs, which produce $\mathcal{O}$(100-1,000) charged particle tracks. Data relations within the collision event can be naturally represented by a heterogeneous graph, with PV and track node types, and track-track and PV-track edges, which signify relations between various node objects. Here, we chose not to include PV-PV edges, as they are not directly associated with any learning objective. Our previous DFEI algorithm assigned each track to a PV with a minimum impact parameter and appended the corresponding PV coordinates as static track‑node features. This capped performance, because the model could not learn to select the correct PV dynamically, and increased memory requirements by duplicating the same coordinates across tracks. By contrast, a heterogeneous graph explicitly encodes PV-track edges, allowing end‑to‑end learning of associations and eliminating repetitions of PV data.
@@
To enable learning of graph representations for joint beauty-hadron reconstruction and PV association, we propose a heterogeneous extension of the GNN layer of Battaglia et al. [45]. We denote the track and PV node representations, $v^{i}_{\rm tr}$ and $v^{j}_{\rm pv}$, and the track-track and PV-track edge representations, $e^{k}_{\rm tr}$ and $e^{l}_{\rm pv-tr}$. The corresponding edge, node and global representation updates for the HGNN layer are as follows:
{split}$\split e^{\prime k}_{\rm tr}=\phi^{e_{\rm tr}}(e^{k}_{\rm tr},v^{r_{k}}_{\rm tr},v^{s_{k}}_{\rm tr},u)\qquad&\bar{e}^{\prime i}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm tr}\})\\ e^{\prime l}_{\rm pv-tr}=\phi^{e_{\rm pv-tr}}(e^{l}_{\rm pv-tr},v^{m_{l}}_{\rm tr},v^{n_{l}}_{\rm pv},u)\qquad\bar{e}^{\prime i}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm pv-tr}\})\\ v^{\prime i}_{\rm tr}=\phi^{v_{\rm tr}}(v^{i}_{\rm tr},\bar{e}^{\prime i}_{\rm tr},\bar{e}^{\prime i}_{\rm pv-tr},u)\qquad\bar{e}^{\prime j}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm pv}}(\{E^{\prime j}_{\rm pv-tr}\})\\ v^{\prime j}_{\rm pv}=\phi^{v_{\rm pv}}(v^{j}_{\rm pv},\bar{e}^{\prime j}_{\rm pv-tr},u)\qquad\bar{e}^{\prime}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow u}(\{E^{\prime}_{\rm tr}\}),\bar{v}^{\prime}_{\rm tr}=\rho^{v_{\rm tr}\rightarrow u}(\{V^{\prime}_{\rm tr}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime}_{\rm tr},\bar{e}^{\prime}_{\rm pv-tr},\bar{v}^{\prime}_{\rm tr},\bar{v}^{\prime}_{\rm pv},u)\qquad\bar{e}^{\prime}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow u}(\{E^{\prime}_{\rm pv-tr}\}),\bar{v}^{\prime}_{\rm pv}=\rho^{v_{\rm pv}\rightarrow u}(\{V^{\prime}_{\rm pv}\})\;,$ (2)
@@
where the HGNN layer update begins with edge updates $e^{\prime k}_{\rm tr}$ and $e^{\prime l}_{\rm pv-tr}$, applying MLPs $\phi^{e_{\rm tr}}$ and $\phi^{e_{\rm pv-tr}}$. Meanwhile, node updates $v^{\prime i}_{\rm tr}$ and $v^{\prime j}_{\rm pv}$ with MLPs $\phi^{v_{\rm tr}}$ and $\phi^{v_{\rm pv}}$ subsequently take aggregated edge representations for multiple edge types as input. The inputs of the $v^{\prime i}_{\rm tr}$ update include the aggregated track edges $\bar{e}^{\prime i}_{\rm tr}$ and PV-track edges $\bar{e}^{\prime i}_{\rm pv-tr}$. Meanwhile, the update $v^{\prime j}_{\rm pv}$ includes the aggregation of PV-track edges to PV nodes, $\bar{e}^{\prime j}_{\rm pv-tr}$. Finally, the global update uses aggregations of all node and edge types as input, which include $\bar{e}^{\prime}_{\rm tr}$, $\bar{e}^{\prime}_{\rm pv-tr}$, $\bar{v}^{\prime}_{\rm tr}$, $\bar{v}^{\prime}_{\rm pv}$. Figure 1(b) shows the various HGNN layer computations of Equation 2 for node and edge inputs from tracks (red) and PVs (blue).
@@
To enhance the scalability of the HGNN layer in potential real-time data acquisition settings, we incorporate edge and node pruning directly into its layers. The goal is to learn node and edge scores that can be used to remove irrelevant nodes and edges at inference time. To this end, we introduce node and edge scores, $\hat{y}^{v}=\sigma(\psi^{v}(v^{\prime}))$ and $\hat{y}^{e}=\sigma(\psi^{e}(e^{\prime}))$, where $\psi^{v}$ and $\psi^{e}$ are MLPs operating on the updated node and edge representations $v^{\prime}$ and $e^{\prime}$ from Equation 2, and $\sigma$ denotes the sigmoid function. These probability scores can be trained in two ways: using a Binary Cross-Entropy loss $\mathcal{L}_{\rm BCE}(\hat{y}^{v/e},y^{v/e})$ when domain knowledge provides suitable ground-truth pruning labels $y^{v}$ and $y^{e}$, or implicitly, by using the predicted scores $\hat{y}^{v}$ and $\hat{y}^{e}$ as weights within the HGNN aggregation functions. We explore both approaches in our ablation studies in Section 4.1.
@@
Pruning can be performed either via a top-$k$ operation or by introducing a threshold $y_{\rm cut}$. In this paper, we adopt the latter approach, given its better observed performance and greater interpretability in terms of purity. Furthermore, we focus on the application of pruning during inference to boost the speed and remove redundant background edges and nodes to enhance reconstruction performance. We acknowledge that integrating pruning into training (i.e., ‘pruning-aware’ training, a form of dynamic graph construction) could further improve the reconstruction. Despite not employing pruning-aware training, our empirical results show that incorporating extra MLPs per node and edge improves the performance of the algorithm, as detailed in the subsequent sections.
@@ 3.3 HGNN architecture and objective loss
The full HGNN architecture is shown in Figure 1(a), which processes the input $G_{\rm in}$ including track and PV node and edge input features through an encoder layer, several HGNN layers with integrated graph pruning and a decoder layer resulting in $G_{\rm out}$. The encoder and decoder layers consist of MLP node, edge and global updates $\chi(v)$, $\chi(e)$ and $\chi(u)$. The encoded graph representation, $G_{\rm enc}$, is concatenated with the output graph representation from each HGNN layer to preserve important information from the input features and prevent oversmoothing [46]. The HGNN was implemented as a custom architecture using the PyTorch framework [47], and the code for training and performance studies is publicly available [48].
The hyperparameters of the HGNN include the number of HGNN layers ($n^{\rm HGNN}_{\rm layers}=8$) and layers within MLPs ($n^{\rm MLP}_{\rm layers}=4$), MLP hidden channel dimensions ($d^{\chi}_{\rm h}=d^{\phi}_{\rm h}=128$ and $d^{\psi}_{\rm h}=16$) and MLP output dimensions ($d^{\chi}_{\rm h}=d^{\phi}_{\rm h}=16$ and $d^{\psi}_{\rm h}=1$). All MLPs employ rectified linear unit (ReLU) activation functions and batch norms, whereas aggregation functions use summation. The hyperparameters were selected based on the observed saturation in the validation loss with model complexity.
The HGNN is trained using stochastic gradient descent with the Adam optimisation routine [49] to accomplish beauty-hadron reconstruction, graph pruning and PV association tasks by minimising a multi-objective loss,
@@
where in the first term $\mathcal{L}_{\rm CE}$ denotes a class weighted cross-entropy loss, $e_{\rm out}$ is the output HGNN edge representation and $y^{\rm LCA}$ is a multi-class target for the LCA reconstruction task. The subsequent terms represent the edge and node pruning tasks introduced in Section 3.2, where the significance of the corresponding pruning task relative to the LCA task is scaled with $\beta$ parameters. The pv-tr edge pruning task effectively performs a PV association by selecting the pv-tr edge with the highest probability score $\hat{y}^{e_{\rm pv-tr}}$ for a given track. By default trainings used $\beta^{e_{\rm tr}}=33$, $\beta^{v_{\rm tr}}=1$ and $\beta^{e_{\rm pv-tr}}=3$. Although a comprehensive optimisation of these hyperparameters was beyond the scope of the study, the sensitivity of the multi-objective optimisation to $\beta^{e_{\rm tr}}$ and $\beta^{v_{\rm tr}}$ is further investigated in  B.
We benchmark the HGNN architecture against a GNN using GNN layer updates according to Equation 1 and an equivalent pruning mechanism. The GNN only considers tracks and their interconnections as nodes and edges. Meanwhile, it uses the same objective loss and hyperparameter values but with the removal of the pv-tr edge pruning task. For both the GNN and HGNN architectures, we also quantify the inclusion of weighted aggregations during message passing and denote the corresponding architectures as WGNN and WHGNN.
@@ 3.4 Datasets and training procedure
The model is trained, validated and tested using the publicly available datasets from our previous publication DFEI [50]. In addition, we provide an updated format of the relevant datasets [51]. The datasets were produced with a custom simulation environment using PYTHIA8 [52] and EvtGen [53] to simulate the particle collision conditions anticipated for LHCb Run 3. Only events containing at least one beauty hadron are considered; these hadrons then decay with EvtGen into a variety of known modes. All results in this paper focus exclusively on charged particles produced within the LHCb geometrical acceptance and Vertex Locator region. Particles outside these regions and neutral particles were not considered, meaning that they are also excluded from the ground truth heavy-hadron decay chains. Additional exclusive datasets, in which one of the beauty hadrons is required to decay to several exclusive decays, are used to further evaluate the model performance.
@@
As depicted in Figure 1, collision events are represented as a heterogeneous graph. Each node and edge type has an input-feature representation. For charged particles (tracks), the input features include the track origin point, $(x_{\rm tr},y_{\rm tr},z_{\rm tr})$, track momenta ($p_{x},p_{y},p_{z}$) and track charge $q$. Meanwhile, PV nodes are represented by their position coordinates, $(x_{\rm pv},y_{\rm pv},z_{\rm pv})$. In contrast, for the GNN benchmarks, we adopted DFEI’s approach of appending the coordinates of the PV with the minimum impact parameter to the features of each track. Although highly discriminating timing information is anticipated for PV association in LHCb Upgrade II, this study limited its scope to using only the previously defined positional and kinematic features.
@@
Initially, a fully connected graph is defined between all track and PV nodes with track-track and PV-track edges. The input edge representation for PV-track edges includes only the impact parameter of the track with respect to the PV. Meanwhile, track-track edge features include the angle $\theta$ between the three-momentum directions of the two particles, the difference $\Delta z_{\rm tr}$ in origin $z_{\rm tr}$, a Boolean indicating a shared PV according to the minimum impact parameter, and the momentum-transverse distance, which is the distance between the origin points in a plane transverse to the momentum direction. We adopted a loose prefiltering of track-track edges with 99% efficiency from the DFEI [22], which requires edges to satisfy $\theta<0.26$ rad. or the condition of a shared PV according to the minimum impact parameter. For the remaining edges, the ground-truth LCAG is determined $y^{\rm LCA}$, where $y^{\rm LCA}=0$ indicates no shared ancestors. The non-zero values were restricted to $y^{\rm LCA}=1,2,3$, which were found to be sufficient to represent the decay hierarchies present in the simulation dataset. Finally, global input features include the multiplicity of tracks and PVs in the event.
@@
The training and validation datasets consisted of 40,000 and 10,000 events respectively, which are equivalent to those used in our previous work. Nominal training for the ablation studies was performed on an NVIDIA L40S GPU with a batch size of 12. For ablation studies, the base schedule consisted of 30 epochs at a learning rate of $\alpha=10^{-3}$, followed by two epochs at $\alpha=10^{-4}$. Before evaluating the reconstruction performance of the selected architectures, we fine-tuned further with two additional epochs at $\alpha=5\times 10^{-5}$ and two epochs at $\alpha=2\times 10^{-5}$. During all training phases, training and validation losses as well as classification accuracies were monitored (see  A for examples of training and validation loss curves). The final performance was measured using a test set of 10,000 independent inclusive beauty-hadron decay events and multiple exclusive decay datasets, each containing 5,000 events.